# Parton Distribution Function of a Deuteron-like Dibaryon System from Lattice QCD

### v1

@@ V Summary
In this paper, We present the first lattice QCD calculation of the quark PDF of the deuteron-like dibaryon system via LaMET method. Our calculation is done on ensembles with a single lattice spacing. With high statistics, we have used multiple source-sink separations to control the excited-state contamination, and applied the state-of-the-art renormalization, matching and extrapolation procedure. We have also studied the momentum dependence, and extrapolated to the infinite momentum limit. For the lattice with a small spatial volume, we find that the PDF exhibits a noticeable difference between the dibaryon system and the sum of a proton and a neutron. As the spatial volume becomes larger, this difference becomes indistinguishable within the current uncertainty.
Clearly, our study requires improvements in various aspects. For example, the current analysis is limited to a single lattice spacing of $0.105$fm, with pion masses of the ensembles at $293$ MeV and $940$ MeV, which are away from the physical point. To assess these systematic uncertainties, a comprehensive analysis using several ensembles with different lattice spacings and pion masses would be essential, which requires a considerable amount of computing resources. Nevertheless, in this work we study, for the first time, the PDFs of a multi-nucleon system. Our final results demonstrate the potential to detect nuclear effects by investigating PDFs through the LQCD approach.
Acknowledgment
@@
We thank the CLQCD collaborations for providing us with their gauge configurations with dynamical fermions Hu et al. (2023), which are generated on HPC Cluster of ITP-CAS, the Southern Nuclear Science Computing Center(SNSC), the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University and the Dongjiang Yuan Intelligent Computing Center. We are grateful to Ying Chen for the valuable discussion and comments. We thank Wei Wang for his constructive feedback on our work . We would like to thank Xiaonu Xiong for providing us a template for plotting the three-point correlator. We would also like to thank Yushan Su for providing the Mathematica notebook to confirm the correct implementation of RGR and LRR. The calculations were performed using the Chroma software suite Edwards and Joo (2005) with QUDA Clark et al. (2010); Babich et al. (2011); Clark et al. (2016) through HIP programming model Bi et al. (2020) and software QCU. The numerical calculations in this paper have been carried out the Dongjiang Yuan Intelligent Computing Center, the HPC Cluster of ITP-CAS and the Southern Nuclear Science Computing Center(SNSC). L. Liu is supported in part by National Natural Science Foundation of China (NSFC) under Grant No. 12293061 and 12175279, the Strategic Priority Research Program of the Chinese Academy of Sciences with Grant No. XDB34030301, Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008. Y. Yang is supported in part by NSFC under Grants No. 12293062, 12293061, 12293065 and 12047503, and also the education integration young faculty project of University of Chinese Academy of Sciences, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. XDB34030303. L. Liu and Y. Yang are supported in part by NSFC under Grants No. 12293060. P. Sun and Y. Yang are supported by Chinese Academy of Sciences under Grant No. YSBR-101. F. Yao and J. H. Zhang are supported in part by National Natural Science Foundation of China under grants No. 12375080, 11975051, and by CUHK-Shenzhen under grant No. UDF01002851. Y. Yang, P. Sun and J. H. Zhang are also supported in part by a NSFC-DFG joint grant under grant No. 12061131006 and SCHA 458/22 and the GHfund A No. 202107011598.
@@ VI.1 Details on data analysis
In Fig. 11, Fig. 12 and Fig. 13 we plotted the ratio of 3pt to 2pt, comparing the fitting results with original data. $z=1a$ results of dibaryon, proton and neutron under different momentum are illustrated in the plots.
Figure 11: Demonstration of fitting the correlation function of ensemble C32P29. Here we represent the result of $z=1a$ of different momentum. We compare the lattice data of ratio (error bars) and results predicted by fitting (colored bands). The ground state contribution $c_{0}$ obtained by the fitting is shown as the black band.
Figure 12: Demonstration of fitting the correlation function of ensemble C24P29. Here we represent the result of $z=1a$ of different momentum.
Figure 13: Demonstration of fitting the correlation function of ensemble C24P90. Here we represent the result of $z=1a$ of different momentum.
In Fig. 14 we present the value of $\chi^{2}$/d.o.f (d.o.f represents degrees of freedom) of correlated fits for $c_{0}$. We take the largest momentum result of dibaryon system depending on different $z$ as an example. Most of the $\chi^{2}$/d.o.f are beneath 1.5 which illustrates the high quality of the fits.
Figure 14: $\chi^{2}$/d.o.f of correlated fits for $c_{0}$, taking the largest momentum of dibaryon system for each ensemble as an example.
@@ VI.2 Renormalization group resummation and leading renormalon resummation
To determine the scheme-dependence $m_{0}$, the method is the comparison between the renormalized matrix elements and the fixed order Wilson coefficient shown in Eq. (20) at a short distance. In principle, we can resum the large logarithmic terms $\alpha_{s}^{n}(\mu)\text{ln}^{n}(z^{2}\mu^{2})$ in $C_{0,\text{NLO}}$ to reduce the scale dependence. Then the fitting parameters $m_{0}$ introduce small deviations from the fixed-order process, changing the z-dependence. Here, the well-known leading infrared renormalon (IRR) arises from the long-distance contributions to the self-energy of the Wilson line, and the asymptotic series is $C(\alpha_{s})=\sum_{i}c_{i}\alpha_{s}^{i}$. The divergence behavior for $c_{i}$ is the same as the quark “pole” mass $m=\sum_{n}r_{n}\alpha_{s}^{n+1}$ that is well investigated Bali et al. (2013). Therefore, the strategy for IRR here is considering LRR proposed in Ref. Zhang et al. (2023) by using state-of-the-art knowledge on the mass renormalon,
@@
×$\times\left(1+\frac{b}{n+b}c_{1}+\frac{b(b-1)}{(n+b)(n+b-1)}c_{2}+...\right)$ (29)
where the parameters b, $c_{1}$, $c_{2}$ and $N_{m}$ are given in Ref. Zhang et al. (2023). One can obtain the leading renormalon contribution through a Borel transformation.
@@
×$\times\frac{1}{(1-2u)^{1+b}}\left(1+c_{1}(1-2u)+c_{2}(1-2u)^{2}+...\right)$ (30)
@@
where the last term $\alpha_{s}(\mu)r_{0}(z,\mu)$ is the NLO expansion of $C_{0,\text{LRR}}^{\text{PV}}(z,\mu)$. Now the result contains the fixed-order contribution and the higher-order leading renormalon contribution. In this work, we resum the leading divergent contributions to all orders at $\mu=z^{-1}$ performing RGR  Su et al. (2023).
In this way, the $m_{0}$ determined using Eq. (VI.2) are shown in Fig. 15 as a function of z. The main values are obtained at $\mu=\sqrt{10}$ GeV by doing global fit. Both statistical and systematic uncertainties are taken into account. The statistical uncertainties remain relatively consistent, regardless of whether LRR and RGR are included. The systematic uncertainties are estimated by varying $\mu$ to $2$ and $4$ GeV. As illustrated in the figure, including the leading renormalon reduces the scale dependence of $m_{0}$. By including LRR and RGR, we improve twist-three power accuracy and high order correction for describing the matrix element in the rest frame, further enhancing the accuracy of the renormalization factor extraction.
Figure 15: $m_{0}$ determined by using fixed order (NLO) and modified (NLO+LRR+RGR) Wilson coefficient as a function of $z$. The main values are obtained at $\mu=\sqrt{10}$ GeV. The bands contain both statistical and systematic uncertainties with the latter determined by varying $\mu$ to 2 and 4 GeV.
@@
In Fig. 16, we present the light-cone PDF of dibaryon system after applying LRR and RGR, with only statistical uncertainties taken into consideration. As can be seen from the figure, the LRR+RGR-improved unpolarized PDF blows up in the small x region , indicating that LaMET is unreliable in this region. In reliable region the NLO+LGR+RGR result shows good agreement with fixed-order matching.
Figure 16: Comparison between fixed-order matching results of dibaryon PDF (blue bands) and that incorporating the LRR+RGR (green bands). The LRR+RGR results blow up at small $x$ region ( $-0.065<x<0.065$), indicating that LaMET is unreliable at this region. Only statistical uncertainties are considered in the plots.

### v2

@@ V Summary
In this paper, we presented the first lattice QCD calculation of the quark PDF of the deuteronlike dibaryon system via the LaMET method. Our calculation was done on ensembles with a single lattice spacing. With high statistics, we have used multiple source-sink separations to control the excited-state contamination and applied the state-of-the-art renormalization, matching, and extrapolation procedure. We have also studied the momentum dependence, and extrapolated the result to the infinite-momentum limit. Finally, we compared the PDF of a deuteronlike dibaryon system with that of free nucleons. The current precision of our calculation was not sufficient enough to draw definitive conclusions on the nuclear effects of dibaryon systems.
In the upcoming EIC and EicC experiments, detecting the nuclear effects related to nuclear PDFs will be one of the most important physical goals. An $ab~{}initio$ calculation of deuteron PDFs will help us understand the related phenomenology. As the first-ever calculation of deuteronlike dibaryon PDFs on lattices, our study requires improvements in various aspects, which will be addressed in future work.
Our analysis was based on a single lattice spacing of $0.105$ fm, which means that discretization errors have not been accounted for. Additionally, the pion masses of the ensembles used were $293$ MeV and $941$ MeV, creating challenges in performing a reliable chiral extrapolation due to the significant gap between these values. Furthermore, at the nonphysical pion mass, it is still an open question whether a bound-state deuteron exists or not on the lattice, complicating the determination of finite-volume effects. Nevertheless, in this work, we studied, for the first time, the PDFs of a multinucleon system. Our final results demonstrated the potential to detect nuclear effects by investigating PDFs through the LQCD approach, with the intention of addressing these challenges in future studies.
ACKNOWLEDGEMENTS
@@
We thank the CLQCD collaborations for providing us with their gauge configurations with dynamical fermions Hu et al. (2023), which are generated on HPC cluster of ITP-CAS, the Southern Nuclear Science Computing Center(SNSC), the Siyuan-1 cluster supported by the Center for High-Performance Computing at Shanghai Jiao Tong University, and the Dongjiang Yuan Intelligent Computing Center. We are grateful to Ying Chen for the valuable discussion and comments. We thank Wei Wang for his constructive feedback on our work. We would like to thank Xiaonu Xiong for providing us a template for plotting the three-point correlator. We would also like to thank Yushan Su for providing the $Mathematica$ notebook to confirm the correct implementation of RGR and LRR. The calculations were performed using the Chroma software suite Edwards and Joo (2005) with QUDA Clark et al. (2010); Babich et al. (2011); Clark et al. (2016) through the HIP programming model Bi et al. (2020) and QCU software. The numerical calculations in this paper were carried out at the Dongjiang Yuan Intelligent Computing Center, the HPC cluster of ITP-CAS and the Southern Nuclear Science Computing Center(SNSC). This work is supported in part by National Natural Science Foundation of China (NSFC) under Grant No. 12293060, No. 12293061, No. 12293062, No. 12175279, No. 12293065, No. 12047503, No. 12435002, No. 12375080 and No. 11975051, the Strategic Priority Research Program of the Chinese Academy of Sciences with Grant No. XDB34030301, No. XDB34030303 and No. YSBR-101, the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008, the Education Integration Young Faculty Project of University of Chinese Academy of Sciences, CUHK-Shenzhen under grant No. UDF01002851, and a NSFC-DFG joint grant under Grant No. 12061131006 and SCHA 458/22 and the GHfund A No. 202107011598.
@@ VI.1 Details on data analysis
In Fig. 13, Fig. 14 and Fig. 15 we plotted the ratio of 3pt to 2pt, comparing the fitting results with original data. $z=1a$ results of dibaryon, proton and neutron under different momenta are illustrated in the plots.
Figure 13: Demonstration of fitting the correlation function of ensemble C32P29. Here we represent the result of $z=1a$ of different momenta. We compare the lattice data of ratios (error bars) and results predicted by fitting (colored bands). The ground-state contribution $c_{0}$ obtained by the fitting is shown as the black band.
Figure 14: Demonstration of fitting the correlation function of ensemble C24P29. Here we represent the result of $z=1a$ of different momenta.
Figure 15: Demonstration of fitting the correlation function of ensemble C24P90. Here we represent the result of $z=1a$ of different momenta.
In Fig. 16 we present the value of $\chi^{2}$/d.o.f. of correlated fits for $c_{0}$. We take the largest-momentum result of the dibaryon system depending on different $z$ as an example. Most of the $\chi^{2}$/d.o.f. are beneath 1.5, which illustrates the high quality of the fits.
Figure 16: $\chi^{2}$/d.o.f. of correlated fits for $c_{0}$, taking the largest momentum of the dibaryon system for each ensemble as an example.
@@ VI.2 Renormalization group resummation and leading renormalon resummation
To determine the scheme dependence $m_{0}$, the method is the comparison between the renormalized matrix elements and the fixed-order Wilson coefficient shown in Eq. (20) at a short distance. In principle, we can resum the large logarithmic terms $\alpha_{s}^{n}(\mu)\text{ln}^{n}(z^{2}\mu^{2})$ in $C_{0,\text{NLO}}$ to reduce the scale dependence. Then the fitting parameters $m_{0}$ introduce small deviations from the fixed-order process, changing the $z$-dependence. Here, the well-known leading infrared renormalon (IRR) arises from the long-distance contributions to the self-energy of the Wilson line, and the asymptotic series is $C(\alpha_{s})=\sum_{i}p_{i}\alpha_{s}^{i}$. The divergence behavior for $c_{i}$ is the same as the quark “pole” mass $m=\sum_{n}r_{n}\alpha_{s}^{n+1}$ that is well investigated Bali et al. (2013). Therefore, the strategy for IRR here is considering LRR proposed in Ref. Zhang et al. (2023) by using state-of-the-art knowledge on the mass renormalon,
@@
×$\times\left(1+\frac{b}{n+b}p_{1}+\frac{b(b-1)}{(n+b)(n+b-1)}p_{2}+...\right)$ (29)
where the parameters $b$, $p_{1}$, $p_{2}$ and $N_{m}$ are given in Ref. Zhang et al. (2023). One can obtain the leading renormalon contribution through a Borel transformation,
@@
×$\times\frac{1}{(1-2u)^{1+b}}\left(1+p_{1}(1-2u)+p_{2}(1-2u)^{2}+...\right)$ (30)
@@
where the last term $\alpha_{s}(\mu)r_{0}(z,\mu)$ is the NLO expansion of $C_{0,\text{LRR}}^{\text{PV}}(z,\mu)$. Now, the result contains the fixed-order contribution and the higher-order leading renormalon contribution. In this work, we resum the leading divergent contributions to all orders at $\mu=z^{-1}$ performing RGR  Su et al. (2023).
In this way, the $m_{0}$ determined using Eq. (VI.2) are shown in Fig. 17 as a function of $z$. The main values are obtained at $\mu=\sqrt{10}$ GeV by doing global fit. Both statistical and systematic uncertainties are taken into account. The statistical uncertainties remain relatively consistent, regardless of whether LRR and RGR are included. The systematic uncertainties are estimated by varying $\mu$ to $2$ and $4$ GeV. As illustrated in the figure, including the leading renormalon reduces the scale dependence of $m_{0}$. By including LRR and RGR, we improve twist-three power accuracy and high-order correction for describing the matrix element in the rest frame, further enhancing the accuracy of the renormalization factor extraction.
Figure 17: $m_{0}$ determined by using fixed-order (NLO) and modified (NLO+LRR+RGR) Wilson coefficient as a function of $z$. The main values are obtained at $\mu=\sqrt{10}$ GeV. The bands contain both statistical and systematic uncertainties with the latter determined by varying $\mu$ to 2 and 4 GeV.
@@
In Fig. 18, we present the light-cone PDF of the dibaryon system after applying LRR and RGR, with only statistical uncertainties taken into consideration. As can be seen from the figure, the LRR+RGR-improved unpolarized PDF blows up in the small x region , indicating that LaMET is unreliable in this region. In the reliable region, the NLO+LGR+RGR result shows good agreement with fixed-order matching.
Figure 18: Comparison between fixed-order matching results of the dibaryon PDF (blue bands) and that incorporating the LRR+RGR (green bands). The LRR+RGR results blow up at the small $x$ region ( $-0.065<x<0.065$), indicating that LaMET is unreliable at this region. Only statistical uncertainties are considered in the plots.